# Topological Corner States in Bilayer and Trilayer Systems with Vertically Stacked Topologically Distinct Layers

### v1

@@ II.2 Topological corner states in BBH trilayer lattice
As we observed a discrepancy between the quadrupole moment and the presence of corner states when $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, in the trilayer case (’topological-trivial-topological’ structure), we need to calculate another topological invariant to discuss its topological phase in detail. We choose the MCN as the additional invariant, which is capable of accurately evaluating the topological phases of $\mathbb{Z}$-class chiral-symmetric higher-order topological systems, as recently introduced by [58]. The MCNs are analogous to the 2D winding number in real space, based on the work described in Ref.[63]. This development is especially crucial in situations where existing theoretical frameworks, such as the nested Wilson loop, may erroneously categorize certain phases as trivial. The MCN provides an innovative approach for analyzing these systems. Notably, the value of MCN directly correlates with the number of degenerate zero-energy corner states present in a finite system, thereby offering a more precise invariant for examining these unique topological states. In this study, we use the MCNs to identify the topological phases of both BBH bilayer and trilayer systems in their chiral-symmetric forms. This is particularly relevant in cases where the conventional nested Wilson approach inaccurately classified the phase as trivial, despite the presence of zero-energy corner states.
To compute the topological invariant for BBH bilayer system, we modify the Hamiltonian in its chiral-symmetric form for its finite-size system as follows:
$\mathcal{H}_{\text{bilayer}}=\begin{pmatrix}0&h_{\text{bilayer}}\\ h^{\dagger}_{\text{bilayer}}&0\end{pmatrix},$ (6)
Note that the labeling of sites in the Hamiltonian differs between layer 1 and layer 2 in order to preserve the chiral symmetry, as shown in Fig. 1(c). Also, sites a and b belong to sublattice A, while sites c and d belong to sublattice B. When a 2D square-lattice has $N$ unit cells in the $x$ and $y$-directions (i.e. each layer consists of $N\times N$ unit cells), $h_{\text{bilayer}}$ is a $4N^{2}\times 4N^{2}$ matrix. Here, the eigenstates of $\mathcal{H}_{\text{bilayer}}$ in equation (6) are denoted as |$\ket{\psi}^{T}=(a_{1}^{1},\cdots,a_{N^{2}}^{1},b_{1}^{1},\cdots,b_{N^{2}}^{1},a_{1}^{2},\cdots,a_{N^{2}}^{2},b_{1}^{2},\cdots,b_{N^{2}}^{2},c_{1}^{1},\cdots,c_{N^{2}}^{1},d_{1}^{1},\cdots,d_{N^{2}}^{1},c_{1}^{2},\cdots,c_{N^{2}}^{2},d_{1}^{2},\cdots,d_{N^{2}}^{2})$. Superscripts represent the layer index, where layer 1 and 2 correspond to upper and lower layer, respectively. The quadrupole moment operator for the BBH bilayer system is represented by a following sublattice quadrupole moment operator[58, 64]
@@
Our calculations reveal that the MCN invariant is $N_{xy}=1$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{bilayer}}$ while $N_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{bilayer}}$, where the MCN number matches the number of corner states at $E=0$. This result is consistent with the Wilson loop calculation. The comprehensive calculations for MCNs are presented in the Appendix D. It is important to note that to calculate the topological invariant, sufficiently large systems are required.
This calculation can readily be adapted for trilayer cases. First, we rewrite the trilayer Hamiltonian in equation (4) and (5) into a form that preserves chiral symmetry, given by
@@
The Hamiltonian $h_{\text{trilayer}}$ is represented by a matrix of size $6N^{2}\times 6N^{2}$. By using the quadrupole moment operators in equation (7), we obtain the MCN, given by equation (8). For a trilayer system depicted in Fig. 1(d), the invariant is obtained as $N_{xy}=2$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$ while $N_{xy}=1$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$. The MCN number perfectly matches the number of corner states at $E=0$. Similarly, for the alternative trilayer system illustrated in Fig. 1(e), we obtain $N_{xy}=1$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$ and $N_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, which again coincides with the number of corner states at $E=0$. Therefore, we conclude that the MCN accurately captures the topological phases that even the quadrupole moment may fail to identify.
@@ IV Discussion and Conclusions
We investigated bilayer and trilayer BBH structures, constructed by stacking topologically distinct layers. By adjusting the interlayer coupling strengths, we identified changes in the number of corner modes, indicating a topological phase transition. This topological phase transition, accompanied by band gap closing, is induced by the interlayer coupling.
@@
To identify the topological phases, we first calculated the quadrupole moments using the nested Wilson loop approach, a well-known method for analyzing quadrupole insulators. The quadrupole moments explained the topological phases of BBH bilayer and ’trivial-topological-trivial’ trilayer structures. However, we observed a mismatch between the quadrupole moments and the presence of corner modes in ’topological-trivial-topological’ trilayer structure when the interlayer coupling was below the critical value. In contrast, MCNs, which are a topological invariant for $\mathbb{Z}$-class chiral-symmetric higher-order topological systems recently reported, can systematically explain the results. We also confirmed that the number of corner states coincides with the MCNs across all interlayer coupling regimes for all three structures we considered. As we identified a topological phase that may have been misclassified as trivial when analyzed through the conventional nested Wilson loop approach, this finding highlights the need for the topological invariant, the MCN, as a more appropriate method of evaluating $\mathbb{Z}$-class higher-order topological systems.
The categorizing the phase transition induced by the interlayer coupling is below:
@@
Figure 6 displays the topological phase diagram, which illustrates both the MCNs and the quadrupole moments in BBH trilayer systems.
This study demonstrates that adjusting the interlayer coupling manipulates the topological phase transition, which is accompanied by the appearance and disappearance of corner modes, potentially expanding the scope of applications. The configuration of vertically stacking layers, as discussed in our study, can be implemented in existing platforms, such as electrical circuits. Additionally, the field distribution of corner modes suggests that the intensity of corner modes can be accumulated through stacked layers, which can be adapted to develop a novel topological laser. We believe that our findings provide valuable insights into the manipulation of localized modes within multilayer systems, opening avenues for innovative device engineering.
@@
Appendix B: Quadrupole moment $q_{xy}$ in BBH system
In the BBH model, the bulk quadrupole moment $q_{xy}$, its boundary polarizations $p_{x,y}$ and corner charges $Q$ have the following relationship: $|p^{\mathrm{edge}\pm y}_{x}|=|p^{\mathrm{edge}\pm x}_{y}|=|Q^{\mathrm{corner}\pm x,\pm y}|=|q_{xy}|$ [27]. The requirements for quantizing the quadrupole moment $q_{xy}$ are existence of three symmetries, which are inversion symmetry and $x,y$-reflection symmetry under a gauge transformation. These constraints lead to quantized values of the Wannier band polarizations $p_{y}^{v_{x}^{\pm}},p_{x}^{v_{y}^{\pm}}\overset{\mathcal{I},\mathcal{M}_{x},\mathcal{M}_{y}}{=}0$ or $1/2$. Furthermore, when the system preserves $C_{4}$ symmetry, the classification of the Wannier bands is $\mathbb{Z}_{2}$ [33]. Therefore, the quadrupole moment $q_{xy}$ can be described by $q_{xy}=p_{y}^{v_{x}^{\pm}}=p_{x}^{v_{y}^{\pm}}=0$ or $1/2$ and the topological invariant is defined by the polarizations which are obtained as follows:
@@
where $\tilde{W}_{y,k_{x}}^{\pm}$ is a nested Wilson loop along $k_{y}$ [27]. The quadrupole moment of $\frac{1}{2}$ and 0 imply that the system is topological and trivial, respectively.
@@
By adapting the Wilson loop calculation to the occupied bands of the BBH bilayer shown in Fig. 2, we obtain four non-degenerate Wannier bands, as depicted in Fig. 7(a1,a2). The gapping of the Wannier bands results from the non-commutativity of the two reflection symmetries, $\mathcal{M}_{x}^{\mathrm{bi}}$ and $\mathcal{M}_{y}^{\mathrm{bi}}$, as discussed in [27]. These two reflection symmetries in the bilayer system are derived from the direct product of the monolayer symmetries, $\mathcal{M}_{x,y}^{\mathrm{mono}}$, expressed as $\mathcal{M}_{x,y}^{\mathrm{bi}}=\mathbb{I}_{2}\otimes\mathcal{M}_{x,y}^{\mathrm{mono}}$.
@@
The Wilson loop calculation applied to the occupied bands of the trilayer structure in Fig. 1(d), yields six non-degenerate Wannier bands (Fig. 7(b1,b2)). As $t_{\mathrm{v}}$ increases, the Wannier bands experience changes, leading to different polarizations. For $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, we obtain polarizations of $p_{y}=\{1/2,1/2,0,0,1/2,1/2\}$, identifying the phase as trivial even with the existence of corner states. In contrast, for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, the polarizations are $p_{y}=\{1/2,0,0,0,0,1/2\}$, indicating a topological phase.
For the alternative trilayer structure depicted in Fig. 1(e), the Wilson loop calculation generates six non-degenerate Wannier bands (Fig. 7(c1,c2)). The nested Wilson loop computation yields polarizations of $p_{y}=\{1/2,0,0,0,0,1/2\}$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$ and $p_{y}=\{0,0,0,0,0,0\}$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$. These polarization changes correspond with the number of corners obtained in finite-sized system calculations.
Appendix C: Perturbation theory
A numerical simulation shows that there are four corner states at large interlayer couplings $t_{\text{v}}$ in the ’topological-trivial-topological’ trilayer structure. In order to understand it, we derive the effective Hamiltonian for large interlayer couplings $t_{\text{v}}$. We first diagonalize the interlayer coupling Hamiltonian $H_{0}$ defined by
@@
$U_{0}^{-1}\mathcal{H}_{0}U_{0}=\left(\begin{array}[]{ccc}-\sqrt{2}\mathbb{I}_{4}&0&0\\ 0&0&0\\ 0&0&\sqrt{2}\mathbb{I}_{4}\end{array}\right),$ (20)
where $\mathbb{I}_{4}$ is the four by four identical matrix and the transformation matrix $U_{0}$ is given by
@@
We are interested in the middle four by four matrix, where there are three degenerate zero energy states in (20). Next, we transform the full Hamiltonian and take the middle four by four matrix, which is given by
@@
where $P_{0}$ is a projection operator taking the middle four by four matrix. It is the effective Hamiltonian, which is also the BBH model. Hence, there are four corner modes for large interlayer couplings $t_{\text{v}}$.
Appendix D: Chiral symmetric Hamiltonian for BBH bilayer system
@@
For $N=2$, the sublattice quadrupole moment operator $\mathcal{Q}_{xy}^{S}$ of the BBH bilayer system is a diagonal matrix where the diagonal elements form a repeating sequence of ($-i,-1,-1,1$), this sequence repeats $2N$ times. Thus, $\mathcal{Q}_{xy}^{S}$ = $\mathbb{I}_{4}\otimes\text{diag.}(-i,-1,-1,1)$.
@@
$h_{\text{trilayer}}=\begin{pmatrix}h_{\text{Topo}}&\mathcal{H}_{\text{vp}}&0\\ \mathcal{H}_{\text{vp}}&h_{\text{Tri}}&\mathcal{H}_{\text{vp}}\\ 0&\mathcal{H}_{\text{vp}}&h_{\text{Topo}}\end{pmatrix},$ (25)

### v2

@@ II.1 Topological corner states in BBH bilayer system
As we observed a discrepancy between the quadrupole moment and the presence of corner states when $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, in the trilayer case (’topological-trivial-topological’ structure), we need to calculate another topological invariant to discuss its topological phase in detail. We choose the MCN as the additional invariant, which is capable of accurately evaluating the topological phases of $\mathbb{Z}$-class chiral-symmetric HOTI systems, as recently introduced by [60]. The MCNs are analogous to the 2D winding number in real space, as established in [69] and recently extended in [70]. This development is especially crucial in situations where existing theoretical frameworks, such as the nested Wilson loop, may erroneously categorize certain phases as trivial. The MCN provides an innovative approach for analyzing these systems. Notably, the value of MCN directly correlates with the number of degenerate zero-energy corner states present in a finite system, thereby offering a more precise invariant for examining these unique topological states. In this study, we use the MCNs to identify the topological phases of both BBH bilayer and trilayer systems in their chiral-symmetric forms. This is particularly relevant in cases where the conventional nested Wilson approach inaccurately classified the phase as trivial, despite the presence of zero-energy corner states.
To compute the topological invariant for the BBH bilayer system, we modify the Hamiltonian in its chiral-symmetric form for its finite-size system as follows:
$\mathcal{H}_{\text{bilayer}}=\begin{pmatrix}0&h_{\text{bilayer}}\\ h^{\dagger}_{\text{bilayer}}&0\end{pmatrix}.$ (6)
Note that the labeling of sites in the Hamiltonian differs between layer 1 and layer 2 in order to preserve the chiral symmetry, as shown in Fig. 1(d). In addition, sites a and b belong to sublattice A, while sites c and d belong to sublattice B. When a 2D square lattice has $N$ unit cells in the $x$ and $y$-directions (i.e., each layer consists of $N\times N$ unit cells), $h_{\text{bilayer}}$ is a $4N^{2}\times 4N^{2}$ matrix. Here, the eigenstates of $\mathcal{H}_{\text{bilayer}}$ in equation (6) are denoted as |$\ket{\psi}^{T}=(a_{1}^{1},\cdots,a_{N^{2}}^{1},b_{1}^{1},\cdots,b_{N^{2}}^{1},a_{1}^{2},\cdots,a_{N^{2}}^{2},b_{1}^{2},\cdots,b_{N^{2}}^{2},c_{1}^{1},\cdots,c_{N^{2}}^{1},d_{1}^{1},\cdots,d_{N^{2}}^{1},c_{1}^{2},\cdots,c_{N^{2}}^{2},d_{1}^{2},\cdots,d_{N^{2}}^{2})$. The superscripts represent the layer index, where layer 1 and 2 correspond to the upper and lower layer, respectively. The quadrupole moment operator for the BBH bilayer system is represented by a following sublattice quadrupole moment operator[60, 71]
@@
Our calculations reveal that the MCN invariant is $N_{xy}=1$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{bilayer}}$ while $N_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{bilayer}}$, where the MCN number matches the number of corner states at $E=0$. This result is consistent with the Wilson loop calculation. The comprehensive calculations for MCNs are presented in Appendix F. It is important to note that, to calculate the topological invariant, sufficiently large systems are required.
This calculation can be readily adopted for trilayer cases. First, we rewrite the trilayer Hamiltonian in equation (4) and (5) into a form that preserves chiral symmetry, given by
@@
The Hamiltonian $h_{\text{trilayer}}$ is represented by a matrix of size $6N^{2}\times 6N^{2}$. By using the quadrupole moment operators in equation (7), we obtain the MCN, given by equation (8). For a trilayer system depicted in Fig. 1(e), the invariant is obtained as $N_{xy}=2$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$ while $N_{xy}=1$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$. The MCN number perfectly matches the number of corner states at $E=0$. Similarly, for the alternative trilayer system illustrated in Fig. 1(f), we obtain $N_{xy}=1$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$ and $N_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, which again coincides with the number of corner states at $E=0$. Therefore, we conclude that the MCN accurately captures the topological phases that even the quadrupole moment may fail to identify.
@@ IV Discussion and Conclusions
We investigated the BBH bilayer and trilayer systems, constructed by stacking topologically distinct layers, forming what we refer to as topological heterostructures. By adjusting the interlayer coupling strengths, we identified changes in the number of corner states, indicating a topological phase transition. This topological phase transition, accompanied by the closing of the band gap, is induced by the interlayer coupling.
@@
To identify the topological phases within these topological heterostructures, we first calculated the quadrupole moments using the nested Wilson loop approach, a well-known method for analyzing quadrupole insulators. The quadrupole moments explained the topological phases of BBH bilayer and ’trivial-topological-trivial’ trilayer structures. However, we observed a mismatch between the quadrupole moments and the presence of corner states in the ’topological-trivial-topological’ trilayer structure when the interlayer coupling was below the critical value. In contrast, MCNs, recently reported as a topological invariant for $\mathbb{Z}$-class chiral-symmetric HOTI systems, can systematically explain the results. We also confirmed that the number of corner states coincides with the MCNs across all interlayer coupling regimes for all three structures we considered.
The categorization of the phase transition induced by interlayer coupling is as follows:
@@
Figure 6 shows the topological phase diagram, which illustrates both the MCNs and the quadrupole moments in BBH trilayer systems.
As this topological phase transition is accompanied by the appearance and disappearance of corner states, these topological heterostructures may offer new possibilities for future applications. The vertically stacked layer configuration discussed in our paper can be implemented using existing platforms, such as electrical circuits. Additionally, the field distribution of the corner modes suggests that their intensity, which extends across multiple layers, can be further enhanced through layer stacking. This finding indicates that topological heterostructures are potentially adoptable for the development of novel topological lasers when implemented on photonic platforms. We believe that our study offers valuable insight into the control of localized corner states in multilayer systems, paving the way for new research directions in vertically stacked topological heterostructures.
@@
Appendix B: The limit model $\kappa_{1}=0$
Corner states that spread over two or three layers exhibit unique distributions in each layer. For example, in the case of a BBH bilayer system composed of topological and trivial, the corner cell within the corner unit cell in layer 1 has an amplitude, while in layer 2, the second cell from the corner exhibits an amplitude. Each layer exhibits staggered finite amplitudes only on a single sublattice of the system, due to chiral symmetric protection [60]. The amplitudes at each site can be explained by analyzing the limit model with $\kappa_{1}=0$.
Figure 7(a) illustrates one corner of the BBH bilayer system in the $\kappa_{1}=0$ limit model. When a single corner mode exists in layer 1, the site amplitudes in layer 2 are given by $a=-b=C_{1}\frac{t_{\text{v}}}{2\kappa_{2}}$, and $c=d=0$.
Figure 7: The limit model $\kappa_{1}=0$ for (a) BBH bilayer, (b1, b2) ’Topological-Trivial-Topological’ BBH trilayer systems and (c) ’Trivial-Topological-Trivial’ BBH trilayer systems. (a, b1, c) Below the critical interlayer coupling strength ($t_{\mathrm{v}}<t_{\mathrm{c}}$). (b2) Above the critical interlayer coupling strength ($t_{\mathrm{v}}>t_{\mathrm{c}}$).
In the $\kappa_{1}=0$ limit model of the BBH trilayer (topological-trivial-topological structure), in-phase corner modes appear in layer 1 and layer 3 for cases below the critical interlayer coupling strength, as shown in Fig. 7(b1). This results in symmetrical coupling, where the site amplitudes in layer 2 are $a=-b=C_{2}\frac{t_{\text{v}}}{\kappa_{2}}$, and $c=d=0$. In contrast, for cases above the critical interlayer coupling strength, out-of-phase corner modes emerge in layer 1 and layer 3, as illustrated in Fig. 7(b2). This results in the absence of amplitude in layer 2 due to anti-symmetric coupling.
In the $\kappa_{1}=0$ limit model of the BBH trilayer (trivial-topological-trivial structure), only a single corner mode exists in layer 2, which belongs to sublattice B. As a result, finite amplitudes appear in the B sublattices of layers 1 and 3, given by $c=-d=C_{3}\frac{t_{\text{v}}}{\kappa_{2}}$, while $a=b=0$. Here, $C_{1},C_{2},C_{3}$ are normalization constants.
Appendix C: Quadrupole moment $q_{xy}$ in the BBH bilayer and trilayer system
In the BBH model, the bulk quadrupole moment $q_{xy}$, its boundary polarizations $p_{x,y}$ and corner charges $Q$ have the following relationship: $|p^{\mathrm{edge}\pm y}_{x}|=|p^{\mathrm{edge}\pm x}_{y}|=|Q^{\mathrm{corner}\pm x,\pm y}|=|q_{xy}|$ [25]. The requirement for quantizing the quadrupole moment $q_{xy}$ is the existence of three symmetries, which are the inversion symmetry and $x,y$-reflection symmetry under a gauge transformation. These constraints lead to quantized values of the Wannier band polarizations $p_{y}^{v_{x}^{\pm}},p_{x}^{v_{y}^{\pm}}\overset{\mathcal{I},\mathcal{M}_{x},\mathcal{M}_{y}}{=}0$ or $1/2$. Furthermore, when the system preserves $C_{4}$ symmetry, the classification of the Wannier bands is $\mathbb{Z}_{2}$ [36]. Therefore, the quadrupole moment $q_{xy}$ can be described by $q_{xy}=p_{y}^{v_{x}^{\pm}}=p_{x}^{v_{y}^{\pm}}=0$ or $1/2$ and the topological invariant is defined by the polarizations which are obtained as follows:
@@
where $\tilde{W}_{y,k_{x}}^{\pm}$ is a nested Wilson loop along $k_{y}$ [25]. The quadrupole moment of $\frac{1}{2}$ and 0 implies that the system is topological and trivial, respectively.
@@
By adopting the Wilson loop calculation to the occupied bands of the BBH bilayer shown in Fig. 2, we obtain four non-degenerate Wannier bands, as depicted in Fig. 8(a1,a2). The gapping of the Wannier bands results from the non-commutativity of the two reflection symmetries, $\mathcal{M}_{x}^{\mathrm{bi}}$ and $\mathcal{M}_{y}^{\mathrm{bi}}$, as discussed in [25]. These two reflection symmetries in the bilayer system are derived from the direct product of the monolayer symmetries, $\mathcal{M}_{x,y}^{\mathrm{mono}}$, expressed as $\mathcal{M}_{x,y}^{\mathrm{bi}}=\mathbb{I}_{2}\otimes\mathcal{M}_{x,y}^{\mathrm{mono}}$.
@@
The Wilson loop calculation applied to the occupied bands of the trilayer structure in Fig. 1(e), yields six non-degenerate Wannier bands (Fig. 8(b1,b2)). As $t_{\mathrm{v}}$ increases, the Wannier bands experience changes, leading to different polarizations. For $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, we obtain the polarizations of $p_{y}=\{1/2,1/2,0,0,1/2,1/2\}$, identifying the phase as trivial even with the existence of corner states. In contrast, for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, the polarizations are $p_{y}=\{1/2,0,0,0,0,1/2\}$, indicating a topological phase.
For the alternative trilayer structure depicted in Fig. 1(f), the Wilson loop calculation generates six non-degenerate Wannier bands (Fig. 8(c1,c2)). The nested Wilson loop computation yields polarizations of $p_{y}=\{1/2,0,0,0,0,1/2\}$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$ and $p_{y}=\{0,0,0,0,0,0\}$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$. These polarization changes correspond to the number of corners obtained in finite-sized system calculations.
Appendix D: Perturbation theory
To provide a clearer understanding of the results obtained from numerical simulations at large coupling strengths $t_{\text{v}}$, we have derived the effective Hamiltonian for these significant interlayer couplings. This derivation reveals both the presence and absence of four corner states within the topological heterostructures under large interlayer couplings.
($\mathrm{i}$) For the bilayer system, we consider the interlayer coupling Hamitonian $H_{0}$, defined by
$\mathcal{H}_{0}=\left(\begin{array}[]{cc}0&\mathcal{H}_{\text{v}}\\ \mathcal{H}_{\text{v}}&0\end{array}\right),$ (19)
where only the interlayer couplings exist. This is diagonalized as
$U_{0}^{-1}\mathcal{H}_{0}U_{0}=\left(\begin{array}[]{cc}-t_{\text{v}}\mathbb{I}_{4}&0\\ 0&t_{\text{v}}\mathbb{I}_{4}\end{array}\right),$ (20)
where $\mathbb{I}_{4}$ is the four-by-four identity matrix. The transformation matrix $U_{0}$ is given by
$U_{0}=\left(\begin{array}[]{cc}-\mathbb{I}_{2}\otimes\sigma_{x}&\mathbb{I}_{4}\\ \mathbb{I}_{2}\otimes\sigma_{x}&\mathbb{I}_{4}\end{array}\right).$ (21)
When extracting either the upper or lower bands, such as extracting the upper band from equation (20) using the projection operator $P_{0}$, where $P_{0}$ projects onto the upper four-by-four submatrix, the resulting Hamiltonian for the upper four-by-four submatrix is given by
$P_{0}^{-1}U_{0}^{-1}\mathcal{H}_{\text{bilayer}}U_{0}P_{0}=\left(\begin{array}[]{cccc}t_{\text{v}}&0&\frac{1}{2}\left(1+e^{ik_{x}}\right)\left(\kappa_{1}+\kappa_{2}\right)&\frac{1}{2}\left(1+e^{ik_{y}}\right)\left(\kappa_{1}+\kappa_{2}\right)\\ 0&t_{\text{v}}&\frac{1}{2}\left(1+e^{-ik_{y}}\right)\left(\kappa_{1}+\kappa_{2}\right)&-\frac{1}{2}\left(1+e^{-ik_{x}}\right)\left(\kappa_{1}+\kappa_{2}\right)\\ \frac{1}{2}\left(1+e^{-ik_{x}}\right)\left(\kappa_{1}+\kappa_{2}\right)&\frac{1}{2}\left(1+e^{ik_{y}}\right)\left(\kappa_{1}+\kappa_{2}\right)&t_{\text{v}}&0\\ \frac{1}{2}\left(1+e^{-ik_{y}}\right)\left(\kappa_{1}+\kappa_{2}\right)&-\frac{1}{2}\left(1+e^{ik_{x}}\right)\left(\kappa_{1}+\kappa_{2}\right)&0&t_{\text{v}}\end{array}\right).$ (22)
At the M-point ($k_{x}=k_{y}=\pi$), all off-diagonal elements in Eq. (22) vanish. Consequently, equation (22) is always gapless at the M-point with an energy offset of $t_{\text{v}}$ independent of the ratio of $\kappa_{1}$ and $\kappa_{2}$. To further investigate the existence of localized states, we calculate the inverse participation ratio (IPR) and confirm the absence of localized corner states at all energy levels in the strong interlayer coupling regime of the bilayer system. The detailed calculations are provided in Appendix E.
($\mathrm{ii}$) For the trilayer ’topological-trivial-topological’ structure, we first diagonalize the interlayer coupling Hamiltonian $H_{0}$ defined by
@@
$U_{0}^{-1}\mathcal{H}_{0}U_{0}=\left(\begin{array}[]{ccc}-\sqrt{2}t_{\text{v}}\mathbb{I}_{4}&0&0\\ 0&0&0\\ 0&0&\sqrt{2}t_{\text{v}}\mathbb{I}_{4}\end{array}\right),$ (24)
and the transformation matrix $U_{0}$ is given by
@@
We are interested in the middle four-by-four matrix, where there are three degenerate zero energy states in equation (24). Next, we transform the full Hamiltonian and take the middle four-by-four matrix, which is given by
@@
where $P_{0}$ is a projection operator taking the middle four-by-four matrix. This results in an effective Hamiltonian, which is equivalent to the BBH model. Consequently, there are four corner modes for large interlayer couplings $t_{\text{v}}$ when $\kappa_{1}<\kappa_{2}$. A similar discussion applies to the absence of corner modes for large interlayer couplings when $\kappa_{1}>\kappa_{2}$, referred to as ’trivial-topological-trivial’, as shown in Fig. 5(a). The detailed calculations are given in Appendix E.
Appendix E: Inverse participation ratio
To investigate the presence of localized states, we calculate the inverse participation ratio (IPR) and confirmed the existence or absence of localized corner states across all energy in BBH bilayer and trilayer systems, both below and above the critical interlayer coupling strength where topological transitions occur.
The IPR quantifies the degree of localization of a wavefunction and is defined as $\text{IPR}=\sum_{i}|\psi_{i}|^{4}$ where $\psi_{i}$ represents the wavefunction at site $i$. A higher IPR value indicates stronger localization, while a lower value corresponds to a more delocalized state.
($\mathrm{i}$) BBH Bilayer System
[…]
($\mathrm{ii}$) BBH trilayer system type $\mathrm{I}$ (topological-trivial-topological structure)
[…]
($\mathrm{iii}$) BBH trilayer system type $\mathrm{II}$ (trivial-topological-trivial structure)
[…]
Figure 9: IPR of (a) BBH bilayer, (b) ’Topological-Trivial-Topological’ BBH trilayer and (c) ’Trivial-Topological-Trivial’ BBH trilayer systems. (a1)-(c1) Below the critical interlayer coupling strength ($t_{\mathrm{v}}=1$). (a2)-(c2) Above the critical interlayer coupling strength ($t_{\mathrm{v}}=6$). Parameters are chosen as $\kappa_{1}=1$, $\kappa_{2}=3$, and $N$=20. The red circle indicates the IPR of the corner states, whereas the green dashed line indicates the ideal IPR value (IPR’) when the corner states are entirely confined and evenly distributed among the corners.
Appendix F: Chiral symmetric Hamiltonian in real space
@@
The sublattice quadrupole moment operator $\mathcal{Q}_{xy}^{S}$ of the BBH bilayer system is a diagonal matrix where the diagonal elements form a repeating sequence of ($-i,-1,-1,1$), this sequence repeats $2N$ times. Thus, $\mathcal{Q}_{xy}^{S}$ = $\mathbb{I}_{4}\otimes\text{diag.}(-i,-1,-1,1)$.
@@
$h_{\text{trilayer}}=\begin{pmatrix}h_{\text{Topo}}&\mathcal{H}_{\text{vp}}&0\\ \mathcal{H}_{\text{vp}}&h_{\text{Tri}}&\mathcal{H}_{\text{vp}}\\ 0&\mathcal{H}_{\text{vp}}&h_{\text{Topo}}\end{pmatrix}.$ (29)